Net income $2,560,000

Sales volume 280,000 units

Selling for $17

VC is $6

FC is $520,000

Management is planning for the next year, where the unit purchase price of the video disks will increase 30%

1. What volume of sales must the company achieve in the coming year to maintain the same net income as projected for the current year if the unit selling price remains at $17?

2. In order to cover a 30% increase in the disks purchase price for the coming year, and still maintain the current contribution margin ratio, what selling price must the company establish for the coming year?
Answer #1

1. Current Contribution margin ratio = (Sales - variable costs)/sales

=(17-6)/17

= 0.6471

Desired contribution margin = (520-000+2560000)=$3,080,000

Margin per unit =(17-6*1.3) =$9.2

Number of units required to be sold = 3080,000/9.2

= 334,782.61 units

2. Required selling price = 6*1.3/(1-0.6471)

=$22.10 per unit